# supplemental_arxiv

### v1

@@ Abstract
A set of probabilities along with corresponding quantiles are often used to define predictive distributions or probabilistic forecasts. These quantile predictions offer easily interpreted uncertainty of an event, and quantiles are generally straightforward to estimate using standard statistical and machine learning methods. However, compared to a distribution defined by a probability density or cumulative distribution function, a set of quantiles has less distributional information. When given estimated quantiles, it may be desirable to estimate a fully defined continuous distribution function. Many researchers do so to make evaluation or ensemble modeling simpler. Most existing methods for fitting a distribution to quantiles lack accurate representation of the inherent uncertainty from quantile estimation or are limited in their applications. In this manuscript, we present a Gaussian process model, the quantile Gaussian process, which is based on established theory of quantile functions and sample quantiles, to construct a probability distribution given estimated quantiles. A Bayesian application of the quantile Gaussian process is evaluated for parameter inference and distribution approximation in simulation studies. The quantile Gaussian process is used to approximate the distributions of quantile forecasts from the 2023-24 US Centers for Disease Control collaborative flu forecasting initiative. The simulation studies and data analysis show that the quantile Gaussian process leads to accurate inference on model parameters, estimation of a continuous distribution, and uncertainty quantification of sample quantiles.
@@ 1 Introduction
The use of quantiles in statistical modeling and in reporting inferential or predictive uncertainty is widespread, and reporting several quantiles or predictive intervals is common for probabilistic forecasting Gneiting et al. (2023). Quantiles can be easier to interpret than statistical model parameters and are often used to define confidence or prediction intervals. Thus with multiple quantiles for a particular outcome, one has a measure of uncertainty. Estimating quantiles in the presence of covariates via quantile regression is a common statistical and machine learning strategy, and where parametric models are complicated or nonexistent quantiles may be easier to estimate via machine learning methods Martin and Syring (2022); Chung et al. (2021); Koenker (2017); Koenker and Bassett Jr (1978). In estimating multiple quantiles, quantile regression often provides a tradeoff with parametric modeling in that for some problems quantile regression may be easier to perform but that predicted quantiles lack the detailed information of a fully defined predictive distribution Pohle (2020). Perhaps, for data privacy reasons, data quantiles are often reported as summaries of the data. Census or medical data, which can be very large or personal to the subjects, may be published as summary or aggregate data including percentiles (quantiles) and medians Simpson et al. (2023); CDC (2022); Nirwan and Bertschinger (2020). In collaborative forecast initiatives and competitions, probabilistic forecasts are often submitted as a set of predictive quantiles Gneiting et al. (2023); Hong et al. (2016). In recent years, disease outbreak forecast hubs require that all forecasts submitted by outside participants be represented by several predictive intervals or quantiles. This standardized representation allows for straightforward forecast scoring and ensemble building Mathis et al. (2024); Github (2024); Cramer et al. (2022a, b); Sherratt et al. (2023); Bracher et al. (2021).
@@
A set of quantiles may provide distributional information for an event, but it is not as informative as a distribution defined by a cumulative distribution function (CDF), the inverse-CDF or quantile function QF, or a probability density/mass function (PDF), and quantiles provide no distribution tail information –information for values below the smallest quantile or above the largest quantile. Another drawback for using quantiles to define a distribution is that many tools for evaluating or scoring continuous distributions require a CDF or PDF Gneiting and Raftery (2007); Gneiting and Katzfuss (2014). Likewise combining distributions into an ensemble distribution is commonly done by aggregating multiple QFs or CDFs/PDFs, the latter being possible only when a CDF/PDF is available Gneiting et al. (2005); Wang et al. (2023).
@@
Fitting quantiles to recover or estimate a continuous distribution is done in many fields, often for the purpose of evaluating forecasts using rules that require a CDF or PDF Simpson et al. (2023); Gerding et al. (2023). Fitting may also be done to allow combining multiple forecasts using aggregation methods that require CDFs or PDFs Gyamerah et al. (2020); Li et al. (2019); Baran and Lerch (2018); Bogner et al. (2017); He et al. (2016); Gneiting et al. (2005). An example of fitting quantiles is given in figure 1. The 12 points in the figure are quantiles estimated from a random sample of size 100 from a standard normal distribution for given probabilities. The quantile function of the standard normal distribution is represented by the solid grey line in the plot. The dashed grey line is the fitted QF of a normal distribution that was estimated by selecting the mean and standard deviation parameters which minimize the least squares distance between the estimated quantiles and the QF.
@@
Hereafter we refer to estimating a continuous distribution by fitting quantiles as quantile matching (QM). Sgouropoulos et al. (2015) performed QM by minimizing the mean square difference between quantiles of a response variable and a linear combination of quantiles of covariates. Selecting distribution parameters which minimize the mean square error between quantiles and a QF is a common QM method Dilger et al. (2022); Li et al. (2019); Belgorodski et al. (2017), kernel density estimation and spline interpolation are common nonparametric QM methods Gerding et al. (2023); Gyamerah et al. (2020); He et al. (2016), and Keelin (2016) introduced a semiparametric method based on defining a flexible QF to fit quantiles. An issue shared by these methods is that they do not adequately account for the uncertainty inherent from the estimation of quantiles. Nirwan and Bertschinger (2020) introduced a QM model based on the definition of sample quantiles as order statistics and formulated a model likelihood for the quantiles as the joint PDF of multiple order statistics. This model provides exact inference for quantile uncertainty, but relies on the CDF and PDF of the distribution, making it less practical for fitting quantile defined distributions which have a QF but for which the CDF and PDF either do not exist in closed form or are difficult to evaluate Perepolkin et al. (2023); Joiner and Rosenblatt (1971); Tukey (1960).
In this manuscript we introduce a novel model, the quantile Gaussian process (QGP), used specifically for QM. The context for using the QGP is that the available data is a set of quantiles estimated for a given set of probabilities. The QGP relies on established theory underlying the relationship between the QF of a continuous distribution and sample quantiles Parzen (2004); Gilchrist (2000); Hyndman and Fan (1996); Walker (1968); Cramér (1951). The QGP can provide accurate asymptotic inference of the quantile uncertainty, and it works well for QM when modeling quantile defined distributions. In section 2 the QF of a continuous random variable and sample quantiles are each defined, and important properties and asymptotic theory are reviewed. In section 3 the QGP model for QM is introduced. Section 4 contains simulation studies illustrating the QGP’s ability to make parameter inference and match the distribution from which the quantiles where estimated. Section 5 contains an application of QM using the QGP model for quantile forecasts of the 2023-24 United States Centers for Disease Control and Prevention (CDC) flu forecasting competition, also known as FluSight. Section 6 concludes this manuscript with a short summary and ideas for further development in QM modeling.
@@ 2 Quantile function and sample quantiles
Along with the CDF and PDF, the QF is a defining function of a random variable. The QF, however, often receives less attention in standard statistical training than do the CDF and PDF Parzen (2004), and the CDF and PDF are more used often for modeling and inference. This section contains the definition of the QF, important properties, the definition of sample quantiles, and key central limit theorems (CLTs) of sample quantiles.
@@ 2.1 Quantile function definition and properties
For most statistical applications, it is useful to define the QF as the inverse of the CDF. However, it is sometimes the case that defining a distribution by a QF rather than a CDF is advantageous. For instance when a distribution is defined by a QF but has no closed form CDF or when simple QFs may be aggregated to define more complex distributions Perepolkin et al. (2023); Gasthaus et al. (2019); Alvarez and Orestes (2023). This section continues with important properties of the QF and functions including the location-scale property, the quantile density function, the probability integral transform, and examples of quantile defined distributions.
@@ 2.1.1 Location-scale property
For a random variable with PDF $f_{0}\left(x\right)$ and QF $Q_{0}\left(p\right)$ for $p\in\left[0,1\right]$, a location-scale family takes the PDF form $\left(1/\sigma\right)f_{0}\left(\left(x-\mu\right)/\sigma\right)$ with location parameter $\mu\in\left(-\infty,\infty\right)$ and scale parameter $\sigma\in\left(0,\infty\right)$ (Casella and Berger, 2002, section 3.5). The QF of the location-scale family then takes the form $\mu+\sigma Q_{0}\left(p\right)$ Parzen (2004). By the addition and multiplication rules from section 3.2 of Gilchrist (2000), this is a valid QF. Gilchrist (2000) outlines the standardization rule which states that if the random variable $Y$ with QF $Q_{Y}\left(p\right)$ has a standard distribution, that is the random variable $Y$ is centered (by a mean or median) at 0 and has a scale of 1, then a random variable with quantile function $Q_{T\left(Y\right)}\left(p\right)=\mu+\sigma Q_{Y}\left(p\right)$ is centered at $\mu$ and has scale $\sigma$.
The linear form for a QF is already inherent in some distributions families such as the normal and logistic distribution families, with respective QFs $\mu+\sigma\Phi^{-1}\left(p\right)$ and $\mu+\sigma\text{log}\left(p/\left(1-p\right)\right)$. For the normal distribution, $\Phi^{-1}\left(p\right)$ is the QF of a standard normal distribution and $\mu$ and $\sigma$ are the mean and standard deviation of the location-scale transformed normal distribution. For the logistic distribution $\text{log}\left(p/\left(1-p\right)\right)$ is the QF of a logistic distribution with mean 0 and scale 1, and $\mu$ and $\sigma$ become the mean and scale parameters of the location-scale transformed logistic distribution. The linear form is convenient for quantile regression modeling and is key to the development of the metalog distribution family which was designed specifically for modeling quantiles Keelin (2016).
@@ 2.1.2 Quantile density function
If $Q(p)$ is the inverse of a CDF $F(x)$, a calculus result shows that the reciprocal QDF $[q(p)]^{-1}=f(Q(p))$, or the PDF evaluated at the QF evaluated at $p$ Perepolkin et al. (2023); Gilchrist (2000). The QDF is important for establishing the CLTs in section 2.2 and for the QGP models introduced in section 3.
@@ 2.1.3 Probability integral transform
An important application of the QF is using it to sample from continuous probability distributions via the probability integral transform (PIT). The PIT states that for any continuous random variable $Y$ with CDF $F_{Y}(y)$, the transformed random variable $X=F_{Y}(Y)$ is uniformly distributed on $(0,1)$, or $X\sim\text{Unif}(0,1)$. Thus if one can sample from a standard uniform distribution, one may also sample from a continuous distribution provided the QF of that distribution can be evaluated or reasonably approximated Wilkinson (2018). The PIT is useful for assessing model fit via generalized residuals Yang (2024); Cox and Snell (1968) and the calibrating of probabilistic forecasts Gneiting et al. (2007). Herein the PIT is important for developing one version of the QGP model, and it is used as part of a distance measure for assessing QM.
@@ 2.1.4 Examples of quantile defined distributions
An example of a quantile defined distribution family that lacks a CDF in closed form is the generalized lambda distribution (GLD) family Ramberg and Schmeiser (1974); Perepolkin et al. (2023). A special case of the GLD family is the Tukey lambda distribution (TLD) family, a one parameter distribution family introduced by Tukey (1960), Joiner and Rosenblatt (1971). The QF for the TLD is in (1) and the QDF in (2).
@@
Another distribution family defined by its quantile function but without a closed form CDF is the metalog distribution (MLD) family, a generalization of the logistic distribution family Keelin (2016). One version of the MLD is defined in (3).
@@ 2.2 Sample quantiles
Using samples of a random variable to estimate quantiles is prevalent in statistical modeling. Given a random sample from some distribution, the sample quantile function may be defined in terms of the order statistics. Hyndman and Fan (1996) review several definitions of sample quantiles used in statistical packages and state definition 2 as a general definition of the functions reviewed.
@@
[…]
Sample quantiles are used to give a probability summary of a dataset. Sample quantiles of Markov chain Monte Carlo draws from a Bayesian posterior distribution are used to analyze posterior distributions, for example by using quantiles to form credible intervals. Well known asymptotic results of sample quantiles are presented below, and it may be noted that asymptotic results for quantile regression quantiles are similar Kocherginsky et al. (2005); Koenker and Bassett Jr (1978).
@@ 2.2.1 Sample quantile central limit theorem
The asymptotic distribution of a set of sample quantiles is given in (5), and it is the basis for the QGP model for QM in section 3. For independent draws $Y_{1},...,Y_{n}$ from a continuous random variable with CDF $F$, PDF $f$, and QF $Q$, and with a set of sample quantiles calculated by (4), theorem 1 holds. Note that in (5), $N_{K}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ refers to a multivariate normal distribution with $K$ dimensions where $\boldsymbol{\mu}$ is the $K$ dimensional mean vector and $\boldsymbol{\Sigma}$ is a $K\times K$ covariance matrix. In (6) the operator $a\wedge b$ is the minimum of the values $a$ and $b$, with order of operations such that $a\wedge b-ab=(a\wedge b)-ab$.
@@
[…]
@@
Theorem 1 has been known since the mid 20th century Cramér (1951), and a thorough though not unique proof is found in Walker (1968). The entries in the covariance matrix in (6) may equivalently be written as
@@
where $q$ is the QDF. Another CLT in (8) for a set of quantiles transformed by the underlying CDF $F(\hat{\boldsymbol{Q}}_{n}(\boldsymbol{p}))=(F(\hat{Q}_{n}(p_{1})),...,F(\hat{Q}_{n}(p_{k})))$ is in (5) and can be derived from (5) by using the PIT and the Delta method Parzen (2004).
[…]
Here the covariance matrix $\Gamma$ has entries $\Gamma_{ij}=p_{i}\wedge p_{j}-p_{i}p_{j}$, making the asymptotic distribution in (8) a Brownian bridge Chow (2009). This second result makes QM using the QGP model possible where $Q(p)$ is difficult to solve as is shown in the following section.

### v2

# Spencer Wadsworth

@@ Abstract
A set of probabilities along with corresponding quantiles are often used to define predictive distributions or probabilistic forecasts. These quantile predictions offer easily interpreted uncertainty of an event, and quantiles are generally straightforward to estimate using standard statistical and machine learning methods. However, compared to a distribution defined by a probability density or cumulative distribution function, a set of quantiles has less distributional information. When given estimated quantiles, it may be desirable to estimate a fully defined continuous distribution function. Many researchers do so to make evaluation or ensemble modeling simpler. Most existing methods for fitting a distribution to quantiles lack accurate representation of the inherent uncertainty from quantile estimation or are limited in their applications. In this manuscript, we present a Gaussian process model, the quantile Gaussian process, which is based on established asymptotic results of quantile functions and sample quantiles, to construct a probability distribution given estimated quantiles. A Bayesian application of the quantile Gaussian process is evaluated for parameter inference and distribution approximation in simulation studies. The quantile Gaussian process is used to approximate the distributions of quantile forecasts from the 2023-24 US Centers for Disease Control collaborative flu forecasting initiative. The simulation studies and data analysis show that the quantile Gaussian process leads to accurate inference on model parameters, estimation of a continuous distribution, and uncertainty quantification of sample quantiles. The method being Bayesian also allows for the use of prior information when approximating a distribution given quantiles.
@@ 1 Introduction
The use of quantiles in statistical modeling and in reporting inferential or predictive uncertainty is widespread, and reporting several quantiles or predictive intervals is common for probabilistic forecasting [1]. Quantiles can be easier to interpret than statistical model parameters and are often used to define confidence or prediction intervals. Thus with multiple quantiles for a particular outcome, one has a measure of uncertainty. Estimating quantiles in the presence of covariates via quantile regression is a common statistical and machine learning strategy, and where parametric models are complicated or nonexistent quantiles may be easier to estimate via machine learning methods [2, 3, 4, 5]. In estimating multiple quantiles, quantile regression often provides a tradeoff with parametric modeling in that for some problems quantile regression may be easier to perform but that predicted quantiles lack the detailed information of a fully defined predictive distribution [6]. Perhaps for data privacy reasons, data quantiles are often reported as summaries of the data. Census or medical data, which can be very large or personal to the subjects, may be published as summary or aggregate data including percentiles (quantiles) and medians [7, 8, 9]. In collaborative forecast initiatives and competitions, probabilistic forecasts are often submitted as a set of predictive quantiles [1, 10]. In recent years, disease outbreak forecast hubs require that all forecasts submitted by outside participants be represented by several predictive intervals or quantiles. This standardized representation allows for straightforward forecast scoring and ensemble building [11, 12, 13, 14, 15, 16].
@@
A set of quantiles may provide distributional information for an event, but it is not as informative as a distribution defined by a cumulative distribution function (CDF), the inverse-CDF or quantile function QF, or a probability density/mass function (PDF), and quantiles provide no distribution tail information –information for values below the smallest quantile or above the largest quantile. Another drawback for using quantiles to define a distribution is that many tools for evaluating or scoring continuous distributions require a CDF or PDF [17, 18]. Likewise combining distributions into an ensemble distribution is commonly done by aggregating multiple QFs or CDFs/PDFs, the latter being possible only when a CDF/PDF is available [19, 20].
@@
Fitting quantiles to recover or estimate a continuous distribution is done in many fields, often for the purpose of evaluating forecasts using rules that require a CDF or PDF [7, 21]. Fitting may also be done to allow combining multiple forecasts using aggregation methods that require CDFs or PDFs [22, 23, 24, 25, 26, 19]. An example of fitting quantiles is given in figure 1. The 12 points in the figure are quantiles estimated from a random sample of size 100 from a standard normal distribution for given probabilities. The quantile function of the standard normal distribution is represented by the solid grey line in the plot. The dashed grey line is the fitted QF of a normal distribution that was estimated by selecting the mean and standard deviation parameters which minimize the least squares distance between the estimated quantiles and the QF.
@@
Hereafter we refer to estimating a continuous distribution by fitting quantiles as quantile matching (QM). [27] performed QM by minimizing the mean square difference between quantiles of a response variable and a linear combination of quantiles of covariates. Selecting distribution parameters which minimize the mean square error between quantiles and a QF is a common QM method [28, 23, 29], kernel density estimation and spline interpolation are common nonparametric QM methods [21, 22, 26], and [30] introduced a semiparametric method based on defining a flexible QF to fit quantiles. An issue shared by these methods is that they do not adequately account for the uncertainty inherent from the estimation of quantiles. [9] introduced a QM model based on the definition of sample quantiles as order statistics and formulated a model likelihood for the quantiles as the joint PDF of multiple order statistics. This model provides exact inference for quantile uncertainty, but relies on the CDF and PDF of the distribution, making it less practical for fitting quantile defined distributions which have a QF but for which the CDF and PDF either do not exist in closed form or are difficult to evaluate [31, 32, 33].
In this manuscript we introduce a novel model, the quantile Gaussian process (QGP) which is used specifically for QM. The context for using the QGP is that the available data is a set of quantiles estimated for a given set of probabilities. The QGP relies on established asymptotic theory underlying the relationship between the QF of a continuous distribution and sample quantiles [34, 35, 36, 37, 38]. The QGP can provide accurate asymptotic inference of the quantile uncertainty, and it works well for QM when modeling quantile defined distributions. The QGP is fit using Bayesian methods and thus allows for using prior information to influence the QM. In section 2 the QF of a continuous random variable and sample quantiles are each defined, and important properties and asymptotic theory are reviewed. In section 3 the QGP model for QM is introduced. Section 4 contains simulation studies illustrating the QGP’s ability to make parameter inference and match the distribution from which the quantiles where estimated. Section 5 contains an application of QM using the QGP model for quantile forecasts of the 2023-24 United States Centers for Disease Control and Prevention (CDC) flu forecasting competition, also known as FluSight. The analyses in sections 4 and 5 show that the QGP accurately estimates distribuions from which the quantiles are estimated and while doing so accurately captures the uncertainty inherent in estimating quantiles. The QGP outperforms all but one other competing methods in estimating the unknown distributions and capturing uncertainty, and also allows for QM in situations where the other methods cannot be used. Section 6 concludes this manuscript with a short summary and ideas for further development in QM modeling.
@@ 2 Quantile function and sample quantiles
Along with the CDF and PDF, the QF is a defining function of a random variable. The QF, however, often receives less attention in standard statistical training than do the CDF and PDF [34], and the CDF and PDF are used more often for modeling and inference. This section contains the definition of the QF, important properties, the definition of sample quantiles, and key central limit theorems (CLTs) of sample quantiles on which the QGP is based.
@@ 2.1 Quantile function definition and properties
For most statistical applications, it is useful to define the QF as the inverse of the CDF. However, it is sometimes the case that defining a distribution by a QF rather than a CDF is advantageous. For instance when a distribution is defined by a QF but has no closed form CDF or when simple QFs may be aggregated to define more complex distributions [31, 39, 40]. This section continues with important properties of the QF and functions including the location-scale property, the quantile density function, the probability integral transform, and examples of quantile defined distributions.
@@ 2.1.1 Location-scale property
For a random variable with PDF $f_{0}\left(x\right)$ and QF $Q_{0}\left(p\right)$ for $p\in\left[0,1\right]$, a location-scale family takes the PDF form $\left(1/\sigma\right)f_{0}\left(\left(x-\mu\right)/\sigma\right)$ with location parameter $\mu\in\left(-\infty,\infty\right)$ and scale parameter $\sigma\in\left(0,\infty\right)$ [41]. The QF of the location-scale family then takes the form $\mu+\sigma Q_{0}\left(p\right)$ [34]. By the addition and multiplication rules from section 3.2 of [35], this is a valid QF. [35] outlines the standardization rule which states that if the random variable $Y$ with QF $Q_{Y}\left(p\right)$ has a standard distribution, that is the random variable $Y$ is centered (by a mean or median) at 0 and has a scale of 1, then a random variable with quantile function $Q_{T\left(Y\right)}\left(p\right)=\mu+\sigma Q_{Y}\left(p\right)$ is centered at $\mu$ and has scale $\sigma$.
The linear form for a QF is already inherent in some distributions families such as the normal and logistic distribution families, with respective QFs $\mu+\sigma\Phi^{-1}\left(p\right)$ and $\mu+\sigma\text{log}\left(p/\left(1-p\right)\right)$. For the normal distribution, $\Phi^{-1}\left(p\right)$ is the QF of a standard normal distribution and $\mu$ and $\sigma$ are the mean and standard deviation of the location-scale transformed normal distribution. For the logistic distribution $\text{log}\left(p/\left(1-p\right)\right)$ is the QF of a logistic distribution with mean 0 and scale 1, and $\mu$ and $\sigma$ become the mean and scale parameters of the location-scale transformed logistic distribution. The linear form is convenient for quantile regression modeling and is key to the development of the metalog distribution family which was designed specifically for modeling quantiles [30].
@@ 2.1.2 Quantile density function
If $Q(p)$ is the inverse of a CDF $F(x)$, a calculus result shows that the reciprocal QDF $q(p)]^{-1}=f(Q(p))$, or the PDF evaluated at the QF evaluated at $p$ [31, 35]. The QDF is important for establishing the CLTs in section 2.2 and for the QGP models introduced in section 3.
@@ 2.1.3 Probability integral transform
An important application of the QF is using it to sample from continuous probability distributions via the probability integral transform (PIT). The PIT states that for any continuous random variable $Y$ with CDF $F_{Y}(y)$, the transformed random variable $X=F_{Y}(Y)$ is uniformly distributed on $(0,1)$, or $X\sim\text{Unif}(0,1)$. Thus if one can sample from a standard uniform distribution, one may also sample from a continuous distribution provided the QF of that distribution can be evaluated or reasonably approximated [42]. The PIT is useful for assessing model fit via generalized residuals [43, 44] and the calibrating of probabilistic forecasts [45]. Herein the PIT is important for developing one version of the QGP model, and it is used as part of a distance measure for assessing QM.
@@ 2.1.4 Examples of quantile defined distributions
An example of a quantile defined distribution family that lacks a CDF in closed form is the generalized lambda distribution (GLD) family [46, 31]. A special case of the GLD family is the Tukey lambda distribution (TLD) family, a one parameter distribution family introduced by [33], [32]. The QF for the TLD is in (1) and the QDF in (2).
@@
Another distribution family defined by its quantile function but without a closed form CDF is the metalog distribution (MLD) family, a generalization of the logistic distribution family [30]. One version of the MLD is defined in (3).
@@ 2.2 Sample quantiles
Using samples of a random variable to estimate quantiles is prevalent in statistical modeling. Given a random sample from some distribution, the sample quantile function may be defined in terms of the order statistics. [36] review several definitions of sample quantiles used in statistical packages and state definition 2 as a general definition of the functions reviewed.
@@
for some $m\in\mathbb{R}$ and $0\leq\gamma\leq 1$ where $\gamma$ is some function of $j=\lfloor pn+m\rfloor$, g=pn+mjg=pn+m-j, and $\lfloor\cdot\rfloor$ is the floor function.
Sample quantiles are used to give a probability summary of a dataset. Sample quantiles of Markov chain Monte Carlo draws from a Bayesian posterior distribution are used to analyze posterior distributions, for example by using quantiles to form credible intervals. Well known asymptotic results of sample quantiles are presented below, and it may be noted that asymptotic results for quantile regression quantiles are similar [47, 5].
@@ 2.2.1 Sample quantile central limit theorem
The asymptotic distribution of a set of sample quantiles is given in (5), and it is the basis for the QGP model for QM in section 3. For independent draws $Y_{1},...,Y_{n}$ from a continuous random variable with CDF $F$, PDF $f$, and QF $Q$, and with a set of sample quantiles calculated by (4), theorem 1 holds. Note that in (5), $N_{K}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ refers to a multivariate normal distribution with $K$ dimensions where $\boldsymbol{\mu}$ is the mean vector and $\boldsymbol{\Sigma}$ is the covariance matrix. In (6) the operator $a\wedge b$ is the minimum of $a$ and $b$, with order of operations such that $a\wedge b-ab=(a\wedge b)-ab$.
@@
$\sqrt{n}(\hat{\boldsymbol{Q}}_{n}(\boldsymbol{p})-\boldsymbol{Q}(\boldsymbol{p}))\overset{D}{\rightarrow}N_{K}(0,\boldsymbol{\mathcal{K}})$ (5)
@@
Theorem 1 has been known since the mid 20th century [38], and a thorough though not unique proof is found in [37]. The entries in the covariance matrix in (6) may equivalently be written as
@@
where $q$ is the QDF. Another CLT in (8) for a set of quantiles transformed by the underlying CDF $F(\hat{\boldsymbol{Q}}_{n}(\boldsymbol{p}))=(F(\hat{Q}_{n}(p_{1})),...,F(\hat{Q}_{n}(p_{k})))$ is in (5) and can be derived from (5) by using the PIT and the Delta method [34].
$\sqrt{n}(F(\hat{\boldsymbol{Q}}_{n}(\boldsymbol{p}))-\boldsymbol{p})\overset{D}{\rightarrow}N_{K}(0,\boldsymbol{\Gamma})$ (8)
Here the covariance matrix $\Gamma$ has entries $\Gamma_{ij}=p_{i}\wedge p_{j}-p_{i}p_{j}$, making the asymptotic distribution in (8) a Brownian bridge [48]. This second result makes QM using the QGP model possible where $Q(p)$ is difficult to solve as is shown in the following section.